# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ 6. Bounding the tree-child SNPR distance
Let $\ell$ be an integer with $\ell>1$. Consider the two tree-child networks ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}^{\prime}_{\ell}$ that are shown in Figure 8. Each of ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}^{\prime}_{\ell}$ has $3\ell$ leaves. Moreover, the agreement tree-child digraph ${\mathcal{D}}_{\ell}$ for ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}^{\prime}_{\ell}$ that is also shown in Figure 8 has cut size $12\ell-$ 1 in each of ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}_{\ell}^{\prime}$. Thus, $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leq 4\ell-2$. We now show that $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=4\ell-2$. Assume that $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})<4\ell-2$. Then there exists an agreement tree-child digraph ${\mathcal{D}}_{\ell}^{*}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ whose cut size in ${\mathcal{N}}_{\ell}$ or ${\mathcal{N}}^{\prime}_{\ell}$ is strictly less than $12\ell-$ 1. Since, for each $j\in\{1,2,\ldots,\ell-1\}$, ${\mathcal{N}}_{\ell}$ displays the two triples $(3j,3j+1,3j+2)$ and $(3j+1,3j+2,3j)$ whereas ${\mathcal{N}}_{\ell}^{\prime}$ only displays the triple $(3j,3j+2,3j+1)$ and no other triple involving $3j$, $13j+$ 1, and $23j+$ 2, a careful check shows that ${\mathcal{D}}^{*}_{\ell}$ contains an element that is not a phylogenetic tree. To see this, note that if ${\mathcal{D}}^{*}_{\ell}$ only consists of phylogenetic trees, then each $j\in\{1,2,\ldots,\ell-1\}$, ${\mathcal{N}}_{\ell}$ contributes two to the cut size of ${\mathcal{D}}^{*}_{\ell}$ in ${\mathcal{N}}_{\ell}$ and two to the cut size of ${\mathcal{D}}^{*}_{\ell}$ in ${\mathcal{N}}_{\ell}^{\prime}$. Thus, there exists an element $D_{i}$ in ${\mathcal{D}}^{*}_{\ell}$ for some $i\in\{\rho,1,2,\ldots,k\}$ that contains a vertex $v$ of in-degree two and out-degree one. Moreover, as ${\mathcal{D}}^{*}_{\ell}$ is an agreement digraph of ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}^{\prime}_{\ell}$, the child of $v$ is $j$ for some $j\in\{4,7,\ldots,3\ell-5\}$. First, assume that there exists a vertex $u$ in $D_{i}$ and two edge-disjoint directed paths $P$ and $P^{\prime}$ from $u$ to $v$. Since ${\mathcal{D}}^{*}_{\ell}$ is tree-child, at least one of $P$ and $P^{\prime}$ contains a vertex $w$ such that the edge $(w,v)$ with $w\neq u$ exists. Furthermore, as $D_{i}$ can be embedded in ${\mathcal{N}}_{\ell}$ and $P$ and $P^{\prime}$ are edge disjoint, we may assume, that the child of $w$ that is not $v$ is $j-1$ or $j+1$. In either case, it is easily seen that there is no embedding of $D_{i}$ in ${\mathcal{N}}_{\ell}^{\prime}$, a contradiction. Thus, we may assume that there exist two vertices $u$ and $u^{\prime}$ with in-degree zero in $D_{i}$ and directed paths from each of $u$ and $u^{\prime}$ to $v$ whose only common vertex is $v$. As $D_{i}$ can be embedded in ${\mathcal{N}}_{\ell}$, we may assume without loss of generality that the child of $u$ is $j+1$ or $j-1$ which leads us to the same contradiction as in the previous case because there is no such embedding of $D_{i}$ in ${\mathcal{N}}^{\prime}_{\ell}$. Hence, there exists no agreement tree-child digraph whose cut size in ${\mathcal{N}}_{\ell}$ or ${\mathcal{N}}^{\prime}_{\ell}$ is strictly less than $12\ell-$ 1. It now follows that
(2) $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=4\ell-2.$
[…]
Figure 8. Two phylogenetic networks ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}_{\ell}^{\prime}$ on $n=3\ell$ leaves with $\ell>1$, an agreement tree-child digraph ${\mathcal{D}}_{\ell}$ for ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}_{\ell}^{\prime}$, an extension of ${\mathcal{D}}_{\ell}$ in ${\mathcal{N}}_{\ell}$ and an extension of ${\mathcal{D}}_{\ell}$ in ${\mathcal{N}}_{\ell}^{\prime}$ indicated by thick lines. This example shows that the bound given in Lemma 6.5 is essentially tight. For details, see the proof of Proposition 6.7.
@@
Turning to $d_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}^{\prime}_{\ell})$, observe that there exists a tree-child SNPR sequence connecting ${\mathcal{N}}_{\ell}$ and ${\mathcal{N}}^{\prime}_{\ell}$ that prunes and regrafts the leaves $21,4,7,\ldots,3\ell-$ 2 in order. Hence,
(3) $d_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}^{\prime}_{\ell})\leq 2\ell.$
By combining Lemma 6.5 with Equations (2) and (3), we have
$2\ell-1=\frac{1}{2}m_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}_{\ell}^{\prime})\leq d_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}^{\prime}_{\ell})\leq 2\ell$
@@
Data availability statement. There is no data associated with the research presented in this paper.
Conflict of interest statement. On behalf of all authors, the corresponding author states that there is no conflict of interest.

### v2

@@ 6 Bounding the tree-child SNPR distance
Let $\ell$ be an integer with $\ell>1$. Consider the two tree-child networks $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$^{\prime}_{\ell}$ that are shown in Figure 14. Each of $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$^{\prime}_{\ell}$ has $3\ell$ leaves. Moreover, the agreement tree-child digraph $\mathcal{D}$$_{\ell}$ for $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$^{\prime}_{\ell}$ that is also shown in Figure 14 has cut size $2\ell-1$ in each of $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$_{\ell}^{\prime}$. Thus, $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leqslant 4\ell-2$. We now show that $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=4\ell-2$. Assume that $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})<4\ell-2$. Then there exists an agreement tree-child digraph ${\mathcal{D}}_{\ell}^{*}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ whose cut size in $\mathcal{N}$$_{\ell}$ or $\mathcal{N}$$^{\prime}_{\ell}$ is strictly less than $2\ell-1$. Since, for each $j\in\{1,2,\ldots,\ell-1\}$, $\mathcal{N}$$_{\ell}$ displays the two triples $(3j,3j+1,3j+2)$ and $(3j+1,3j+2,3j)$ whereas $\mathcal{N}$$_{\ell}^{\prime}$ only displays the triple $(3j,3j+2,3j+1)$ and no other triple involving $3j$, $3j+1$, and $3j+2$, a careful check shows that $\mathcal{D}$$^{*}_{\ell}$ contains an element that is not a phylogenetic tree. To see this, note that if $\mathcal{D}$$^{*}_{\ell}$ only consists of phylogenetic trees, then each $j\in\{1,2,\ldots,\ell-1\}$, $\mathcal{N}$$_{\ell}$ contributes two to the cut size of $\mathcal{D}$$^{*}_{\ell}$ in $\mathcal{N}$$_{\ell}$ and two to the cut size of $\mathcal{D}$$^{*}_{\ell}$ in $\mathcal{N}$$_{\ell}^{\prime}$. Thus, there exists an element $D_{i}$ in $\mathcal{D}$$^{*}_{\ell}$ for some $i\in\{\rho,1,2,\ldots,k\}$ that contains a vertex $v$ of in-degree two and out-degree one. Moreover, as $\mathcal{D}$$^{*}_{\ell}$ is an agreement digraph of $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$^{\prime}_{\ell}$, the child of $v$ is $j$ for some $j\in\{4,7,\ldots,3\ell-5\}$. First, assume that there exists a vertex $u$ in $D_{i}$ and two edge-disjoint directed paths $P$ and $P^{\prime}$ from $u$ to $v$. Since $\mathcal{D}$$^{*}_{\ell}$ is tree-child, at least one of $P$ and $P^{\prime}$ contains a vertex $w$ such that the edge $(w,v)$ with $w\neq u$ exists. Furthermore, as $D_{i}$ can be embedded in $\mathcal{N}$$_{\ell}$ and $P$ and $P^{\prime}$ are edge disjoint, we may assume, that the child of $w$ that is not $v$ is j1j-1 or $j+1$. In either case, it is easily seen that there is no embedding of $D_{i}$ in $\mathcal{N}$$_{\ell}^{\prime}$, a contradiction. Thus, we may assume that there exist two vertices $u$ and $u^{\prime}$ with in-degree zero in $D_{i}$ and directed paths from each of $u$ and $u^{\prime}$ to $v$ whose only common vertex is $v$. As $D_{i}$ can be embedded in $\mathcal{N}$$_{\ell}$, we may assume without loss of generality that the child of $u$ is $j+1$ or j1j-1 which leads us to the same contradiction as in the previous case because there is no such embedding of $D_{i}$ in $\mathcal{N}$$^{\prime}_{\ell}$. Hence, there exists no agreement tree-child digraph whose cut size in $\mathcal{N}$$_{\ell}$ or $\mathcal{N}$$^{\prime}_{\ell}$ is strictly less than $2\ell-1$. It now follows that
$m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=4\ell-2.$ (2)
@@
Turning to $d_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}^{\prime}_{\ell})$, observe that there exists a tree-child SNPR sequence connecting $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$^{\prime}_{\ell}$ that prunes and regrafts the leaves $1,4,7,\ldots,3\ell-2$ in order. Hence,
$d_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}^{\prime}_{\ell})\leqslant 2\ell.$ (3)
By combining Lemma 14 with Equations (2) and (3), we have
$2\ell-1=\frac{1}{2}m_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}_{\ell}^{\prime})\leqslant d_{\rm tc}({\mathcal{N}}_{\ell},{\mathcal{N}}^{\prime}_{\ell})\leqslant 2\ell$
@@
7 Concluding remarks
In this paper, we have taken a step towards approximating the tree-child SNPR distance $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ between two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. By introducing phylogenetic digraphs and their extensions, thereby generalising agreement forests for two phylogenetic trees to two phylogenetic networks, we have shown that $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ is tightly bounded from above and below within small constant factors of $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$. A possible next step is the development of an algorithm to compute $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$. Due to the intricacies of phylogenetic digraphs and their embeddings, this is a major challenge. In addition, it immediately follows from Proposition 15 and the NP-hardness of computing the rSPR distance between two phylogenetic trees $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ [6] that computing $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ is also NP-hard. Since it seems natural to assume that any algorithm for computing $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ needs to repeatedly compute cut sizes, it would be interesting to investigate if the cut size of a given phylogenetic digraph for a phylogenetic network $\mathcal{N}$} can be computed efficiently. In a different direction, the development of reductions and divide-and-conquer strategies for computing $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ is another avenue for future research. For example, in the context of phylogenetic trees, the introduction of the subtree and chain reduction has led to fixed-parameter tractable algorithms for computing the rSPR distance [6]. Do similar reductions exist for phylogenetic networks?
As mentioned in the introduction, we use two different weights for SNPR operations to make our approach work. Specifically, each SNPR+ and SNPR- is weighted one and each SNPR± is weighted two. While these weights differ from the uniform weights that are used for computing the rSPR distance between two phylogenetic trees, they are a consequence of how our generalisation from agreement forests to agreement digraphs and their embeddings works. Without going into detail, given two phylogenetic trees $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ with $d_{\mathrm{rSPR}}({\mathcal{T}},{\mathcal{T}}^{\prime})=k$, there exists an agreement forest $\mathcal{F}$} for $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ with $k+1$ components. Moreover, one can obtain $\mathcal{F}$} from $\mathcal{T}$} (resp. $\mathcal{T}$$^{\prime}$) by deleting $k$ edges in $\mathcal{T}$} (resp. $\mathcal{T}$$^{\prime}$) and suppressing vertices with in-degree one and out-degree one after each edge deletion. Intuitively, each rSPR operation is witnessed by an edge in $\mathcal{T}$} and by an edge in $\mathcal{T}$$^{\prime}$. In the language of this paper, any agreement forest $\mathcal{F}$} for $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ has the property that its cut size in $\mathcal{T}$} is equal to its cut size in $\mathcal{T}$$^{\prime}$ and, thus, $d_{\mathrm{rSPR}}({\mathcal{T}},{\mathcal{T}}^{\prime})$ simply equates to the cut size of $\mathcal{F}$} in one of the two trees. Now consider two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ such that $\mathcal{N}$$^{\prime}$ can be obtained from $\mathcal{N}$} by a sequence $\sigma$ of only SNPR+ operations. Suppose that the length of $\sigma$ is $k$. In this case, $\mathcal{N}$} is an agreement tree-child digraph $\mathcal{D}$} for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, the cut size of $\mathcal{D}$} in $\mathcal{N}$} is zero, and the cut size of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ is $k$. More generally, for an arbitrary tree-child SNPR sequence that connects two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, each SNPR± contributes to the cut sizes of both trees, whereas each SNPR+ and SNPR- only contributes to the cut size of one tree. Thus any approach for computing $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ that is based on cut sizes as defined in this paper (probably) needs to apply non-uniform weights to the different types of SNPR operations. Ultimately, it would be interesting to investigate whether or not an approach exists for computing $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ that allows for any combination of weights.
Acknowledgements
We thank the two anonymous referees for their careful reading and constructive comments. All authors thank the New Zealand Marsden Fund for their financial support. Part of this paper is based upon work supported by the National Science Foundation under Grant No. DMS-1929284 while the second and third authors were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, Rhode Island, US, during the Theory, Methods, and Applications of Quantitative Phylogenomics program.